# Graph Neural Network Surrogates to leverage Mechanistic Expert Knowledge towards Reliable and Immediate Pandemic Response

### v3

@@ 3.2 Neural network performance without spatial resolution
We consider simple neural networks, such as multilayer perceptrons (MLPs), convolutional neural networks (CNNs) with one-dimensional convolutions in the time dimension, and long short-term memory (LSTM) networks. We conduct grid searches on the number of hidden layers (0 to 4) and neurons (neurons, channels, or units) per layer (32 to 1024) for a setting with six age groups and simulated over 30 days without contact change points. We used an 80-10-10 split of the dataset into training, validation, and test sets. All grid searches are conducted by using a five-fold cross validation (CV) to avoid artifacts from the data split. The logarithmically scaled MAPE values of the cross-validation are shown in Fig. 4 (Panel A). The MAPE on the validation sets will be denoted as validation MAPE, the MAPE on a corresponding test set will be denoted as test MAPE. A further analysis revealed Adam as the best optimizer and ELU as the best activation function; see Fig. 4 (Panel B). While we evaluated the models on both datasets, outbreak and persistent-threat, we only provide the latter case, which is the substantially more complex one. The best architecture in our comparison was a simple LSTM with one hidden layer and 1024 units which achieved a MAPE of 0.49 % across the grid search and, thus, 3-4 magnitudes better than our dummy baseline estimator. We further test it on longer time horizons with up to three contact change points; see Fig. 4 (Panel C). While the architecture is optimized on a much simpler model, we obtain test MAPEs between only 5-10 % for the input features depicted in Fig. 3 (Panel A). In the supplementary material, Figs. 7 and 8, we provide predictions for the first two trajectories with and without stratification into age groups to visually assess the meaning of the corresponding MAPE values.
@@ 3.3 Optimized GNN architecture
Combining the results from our explorational experiments of the simple MLP and LSTM neural networks with some minor adaptations, we use the input data set structure for the GNNs as presented in the previous section and shown in Fig. 3 (Panel B). For a GNN to be trained or executed, we provide an (additional) adjacency matrix $A$ for the considered regions. The adjacency matrix is naturally obtained from the graph-based expert modeling approach. The dimension of the adjacency matrix is $400\times 400$, which comes from the particular application on the 400 German counties. Similarly, the first array dimension of the remaining input is 400. Eventually, to adequately integrate heterogeneous disease dynamics over an entire country, we randomize historic and ongoing disease dynamics over all 400 counties. The ongoing disease dynamics are initialized per county with the persistent-threat scenario, thus allowing between very small and very high infectious disease dynamics. For the first four initializations, see Fig. 5 (Panel A). To additionally randomize for the historic dynamics and immunities, we draw the number of recoveries uniformly from the interval between zero and one minus the maximum ongoing dynamics (infected and dead) that are possible in the persistent-threat range.
@@
Having derived a suitable set of input features for a GNN surrogate model, we selected an appropriate GNN model and optimized its architecture and hyperparameters. We conducted a grid search to identify the model architecture with the best performance, also taking into account the respective training time, which can be considerable. To avoid unnecessarily long training times, we used an early stopping criterion after 200 iterations without improvement (i.e., patience of 200 iterations). We start by considering different graph layers as described in the Materials and Methods section and vary the number of layers between 1 and 3 and the number of channels per layer from 32 to 1024.
@@
We observe that the ARMAConv model with three layers and 512 channels per layer performs best with respect to the validation MAPE (Fig. 5, Panel B), while only the GATConv models with the highest MAPEs need substantially less training time (Fig. 5, Panel C). We thus refine the grid search for the ARMAConv model to seven layers and observe that the models with six or seven layers and 512 channels per layer perform best. While the validation MAPE for seven layers, 9.03 %, is slightly larger than that for six layers, 8.98 % (Fig. 5, Panel D), the training time for seven layers, 31.46 minutes, is reduced by almost a factor of two compared with the model with six layers, which requires 58.38 minutes of training (Fig. 5, Panel E). In Fig. 5 (Panel F), we report an additional assessment on hyperparameter tuning, testing the parameters that are specific to the ARMAConv model: the numbers of parallel GCS stacks and iterations per stack. For details, we again refer to the Materials and Methods section and the original publication of the ARMAConv model [49].
@@
Figure 5: Spatially distributed disease dynamics and grid search of different GNN infrastructures. A) Four random initializations of heterogeneous disease dynamics over the 400 counties of Germany. B) Validation MAPEs for the four different GNN layers. C) Training times corresponding to Panel B). D) Extended grid search for ARMAConv model. E) Training times corresponding to Panel D). F) Further hyperparameter tuning for ARMAConv model with seven layers and 512 channels per layer.
@@
Figure 6: Test MAPE for final ARMAConv GNN and runtime comparison between GNN surrogate model and expert mechanistic model. A) Test MAPE over the 400 counties. B) Test MAPE over prediction horizons of 30, 60, and 90 days without contact change and with 0, 1, 2, and 3 contact changes over 30 days horizon. C) Runtime comparison between final GNN and mechanistic expert model of [27] for one, ten, and 100 model executions with either zero or three contact changes over different prediction horizons.
@@ 3.4 Execution Time Performance
While the time and energy needed for training a surrogate model should also be taken into account, the execution time performance is the critical factor if the model is to be used with a graphical user interface (GUI) that needs to respond immediately. The best GNN surrogate model from the previous section yielded test MAPEs between approximately 10 % and 27 %, when predicting spatially and demographically resolved disease dynamics over the 400 counties of Germany. In this section, we present result that indicate how the surrogate model outperforms the expert mechanistic model in terms of execution times. Timings were measured on a GPU-accelerated Intel Xeon Scalable Processor “Skylake” Silver node with four NVIDIA Tesla V100 SXM2 GPUs with 16 GB, utilizing a single GPU for the measurements.
@@ 4 Discussion
Surrogate models are a cost-efficient technique to approximate the behavior of complex models. Usually, the more complex the reference model is, the higher the requirement of computational resources. The goal of the surrogate model is to approximate the original model’s behavior, achieving accuracies close to the original model, at substantially reduced computational cost and time. In time-critical epidemic or pandemic situations, the evaluation of a large set of potential reactions is recommended, and swift reaction is crucial [52]. Unfortunately, supercomputing resources might not be available on model execution or predictions might be needed by stakeholders which do not have access to supercomputing facilities. To overcome these barriers and provide just-in-time answers to urgent questions, (pre)trained, e.g., overnight, surrogate models can become of great aid for decision makers and substantially increase the impact of mathematical-epidemiological modeling. Furthermore, most parameters are already preselected, which simplifies usage and lowers the entry barrier, as users require less prior knowledge to operate the black-box model.
In this article, we present a surrogate model based on neural networks that is capable of handling temporal and spatial resolution. While several recent surrogate models are based on the concept of physics-informed neural networks [53, 54, 55], our surrogate model is entirely data-driven and uses a “physics”-based simulator only to create a wide range of reliable training data. On the one hand, this means that physical constraints such as population conservation are not considered explicitly. On the other hand, our approach offers a substantial advantage, as particular modifications to the original model do not require changes in the surrogate model architecture; only new training data has to be generated and the model has to be retrained, potentially using the previous model as a starting point. Since (epidemiological) parameters usually do not change on a daily basis, retraining is only necessary if these parameters change substantially. In future work, we will investigate if our surrogate model could further benefit from integrating “physical” constraints such as population conservation into the loss function.
@@
In this study, we first compare four different types of GNN layers using the binary adjacency matrix modeling the mobility between different spatial regions as input. Therefore, although already providing convincing results, none of the considered approaches takes the highly varying values for traveling and commuting activities between the different spatial regions into account. We expect that incorporating this information as edge weights in the input graph could further improve the performance. Promising directions for incorporating the edge weights could be the CrystalConv [56] or XENetConv [57] layers. However, due to hardware limitations, specifically insufficient GPU memory on a single NVIDIA Quadro RTX 6000, we were unable to train models incorporating large graphs with 40,000 edges. Addressing this limitation will be a focus of future work.
Eventually, this study sheds light on the performance of a GNN surrogate model using the same graph connectivity in training and evaluation processes. Here, we did not explore the generalization potential of the GNN surrogate model to other graphs than the one used for training, which is one of the strengths of GNNs. We did not consider this as we are focusing solely on the 400 German counties. However, exploring the generalization, incorporating also edge weights, would be an interesting avenue for future research.
@@
Our surrogate model delivers spatially and demographically resolved disease dynamics for a prediction horizon of several months and speeds up the prediction process substantially. As these results can be computed within a fraction of a second, our model is capable of being integrated in web-based application.
@@ Acknowledgements
All data and code used to generate the results presented in this work are publicly available as part of the MEmilio framework [58]. The specific implementation used in this study, including all relevant scripts and pretrained models, can be found under a persistent doi at zenodo [59].
Appendix A Supplementary Figures

### v4

@@ 3.2 Neural network performance without spatial resolution
We consider simple neural networks, such as multilayer perceptrons (MLPs), convolutional neural networks (CNNs) with one-dimensional convolutions in the time dimension, and long short-term memory (LSTM) networks. We conduct grid searches on the number of hidden layers (0 to 4) and neurons (neurons, channels, or units) per layer (32 to 1024) for a setting with six age groups and simulated over 30 days without contact change points. We used an 80-10-10 split of the dataset into training, validation, and test sets. All grid searches are conducted by using a five-fold cross validation (CV) to avoid artifacts from the data split. The logarithmically scaled MAPE values of the cross-validation are shown in Fig. 5 (Panel A). The MAPE on the validation sets will be denoted as validation MAPE, the MAPE on a corresponding test set will be denoted as test MAPE. A further analysis revealed Adam as the best optimizer and ReLU as the best activation function; see Fig. 5 (Panel B). While we evaluated the models on both datasets, outbreak and persistent-threat, we only provide the latter case, which is the substantially more complex one. The best architecture in our comparison was a simple LSTM with one hidden layer and 1024 units which achieved a MAPE of 0.49 % across the grid search and, thus, 3-4 magnitudes better than our dummy baseline estimator. We further test it on longer time horizons with up to three contact change points; see Fig. 5 (Panel C). While the architecture is optimized on a much simpler model, we obtain test MAPEs between only 5-10 % for the input features depicted in Fig. 4 (Panel A). In the supplementary material, Figs. 10 and 11, we provide predictions for the first two trajectories with and without stratification into age groups to visually assess the meaning of the corresponding MAPE values.
@@
Figure 6: Spatially distributed disease dynamics and grid search of different GNN infrastructures. A) Four random initializations of heterogeneous disease dynamics over the 400 counties of Germany. B) Validation MAPEs (in log-scale) for the four different GNN layers. C) Training times corresponding to Panel B). D) Extended grid search for ARMAConv model (in log-scale). E) Training times corresponding to Panel D). F) Further hyperparameter tuning for ARMAConv model with seven layers and 512 channels per layer (in log-scale).
@@ 3.3 Optimized GNN architecture
Combining the results from our explorational experiments of the simple MLP and LSTM neural networks with some minor adaptations, i.e., eliminating the redundant storage of the contact matrices of the 2D dimensional $5\times 162$ and transforming it to a one dimensional vector of length 354, we use the input data set structure for the GNNs as presented in the previous section and shown in Fig. 4 (Panel B). For a GNN to be trained or executed, we provide an (additional) adjacency matrix $A$ for the considered regions. The adjacency matrix is naturally obtained from the graph-based expert modeling approach. The dimension of the adjacency matrix is $400\times 400$, which comes from the particular application on the 400 German counties. Similarly, the first array dimension of the remaining input is 400. Eventually, to adequately integrate heterogeneous disease dynamics over an entire country, we randomize historic and ongoing disease dynamics over all 400 counties. The ongoing disease dynamics are initialized per county with the persistent-threat scenario, thus allowing between very small and very high infectious disease dynamics. For the first four initializations, see Fig. 6 (Panel A). To additionally randomize for the historic dynamics and immunities, we draw the number of recoveries uniformly from the interval between zero and one minus the maximum ongoing dynamics (infected and dead) that are possible in the persistent-threat range.
@@
Having derived a suitable set of input features for a GNN surrogate model, we selected an appropriate GNN model and optimized its architecture and hyperparameters. We conducted a grid search to identify the model architecture with the best performance, also taking into account the respective training time, which can be considerable. To avoid unnecessarily long training times, we used an early stopping criterion after 200 iterations without improvement (i.e., patience of 200 iterations). Again, averaged validation MAPEs are presented in log-scale.
@@
We start by considering different graph layers as described in the Materials and Methods section and vary the number of layers between 1 and 3 and the number of channels per layer from 32 to 1024. We observe that the ARMAConv model with three layers and 512 channels per layer performs best with respect to the validation MAPE (Fig. 6, Panel B), while only the GATConv models with the highest MAPEs need substantially less training time (Fig. 6, Panel C). We thus refine the grid search for the ARMAConv model to seven layers and observe that the models with six or seven layers and 512 channels per layer perform best. While the validation MAPE for seven layers, 9.03 %, is slightly larger than that for six layers, 8.98 % (Fig. 6, Panel D), the training time for seven layers, 31.46 minutes, is reduced by almost a factor of two compared with the model with six layers, which requires 58.38 minutes of training (Fig. 6, Panel E). This counterintuitive reduction in total training time results from the early stopping mechanism: although the computational cost per epoch increases with model depth, the 7-layer model converged in substantially fewer epochs across the cross-validation folds compared to the 6-layer model. To ensure that this large architecture remains stable, we analyze the learning dynamics. As shown in Fig. 7, training and validation errors decrease without divergence. At the best epoch (1277), the validation MAPE is comparable to the training MAPE, confirming robust generalization without overfitting.
Figure 7: Learning dynamics of the optimized ARMAConv model (7 layers, 512 channels). The training (red) and validation (green) MAPEs show consistent convergence.
In Fig. 6 (Panel F), we report an additional assessment on hyperparameter tuning, testing the parameters that are specific to the ARMAConv model: the numbers of parallel GCS stacks and iterations per stack. For details, we again refer to the Materials and Methods section and the original publication of the ARMAConv model bianchi_graph_2021 .
@@
Figure 8: Generalization of GNN surrogate under edge removal. The bar chart shows MAPE (original scale) for four scenarios: baseline, 10% edge removal, 20% edge removal, and 40% edge removal. Bars represent mean prediction error across 20 independent epidemic simulations per scenario using the pre-trained model without retraining. The red dashed horizontal line marks baseline performance on the original 400-county German mobility network.
3.4 Generalization to modified networks
In this section, we briefly report on the generalization capabilities of the trained GNN. We use the model which was trained on a fixed graph topology where all $1,000$ training samples used the same mobility network.
We then created three modified mobility scenarios by randomly removing edges from the original network with (b) 10 % edge removal, (c) 20 % edge removal, and (d) 40 % edge removal, always comparing against the baseline (0% reduction).
For each scenario, we generated 20 independent simulations using the mechanistic ODE metapopulation model with the modified mobility matrices using the persistent threat scenario. We then evaluated the pre-trained GNN model on these novel test sets and measured prediction accuracy using MAPE on the original scale. The results are presented in Fig. 8 and demonstrate generalization capabilities for graph perturbations with 10 or 20 % less edges and clear degradation as topological changes become substantial. With 10 % edge removal, MAPE increases from 8.95% to 14.37%. The model maintains reasonable prediction accuracy, suggesting robustness to minor network disruptions (e.g., temporary mobility restrictions in specific regions). With 20 % edge removal, MAPE rises to 23.01%. While noticeable, predictions may still provide useful scenario guidance with the GNN showing sensitivity to moderate topological changes. With 40 % edge removal, MAPE jumps to 176.56%, indicating high prediction errors and when nearly half of the mobility network is removed.
In summary, the surrogate is reasonably robust to modest mobility changes (e.g., 10 to 20 % edge reduction for simulating localized travel restrictions) but sensitive to major structural changes of the mobility network. On the other side, the mechanistic model itself is robust to any topology change, as it simulates dynamics from first principles and the surrogate trades this flexibility for computational speed.
Figure 9: Test MAPE for final ARMAConv GNN and runtime comparison between GNN surrogate model and expert mechanistic model. A) Test MAPE over the 400 counties. B) Test MAPE over prediction horizons of 30, 60, and 90 days without contact change and with 0, 1, 2, and 3 contact changes over 30 days horizon. C) Runtime comparison between final GNN and mechanistic expert model of kuhn_assessment_2021 for one, ten, and 100 model executions with either zero or three contact changes over different prediction horizons.
@@ 3.5 Execution Time Performance
While the time and energy needed for training a surrogate model should also be taken into account, the execution time performance is the critical factor if the model is to be used with a graphical user interface (GUI) that needs to respond immediately. The best GNN surrogate model from the previous section yielded test MAPEs between approximately 10 % and 27 %, when predicting age-structured and spatially resolved disease dynamics over the 400 counties of Germany. In this section, we present result that indicate how the surrogate model outperforms the expert mechanistic model in terms of execution times. Timings were measured on a GPU-accelerated Intel Xeon Scalable Processor “Skylake” Silver node with four NVIDIA Tesla V100 SXM2 GPUs with 16 GB, utilizing a single GPU for the measurements.
@@ 4 Discussion
Surrogate models are a cost-efficient technique to approximate the behavior of complex models. Usually, the more complex the reference model is, the higher the requirement of computational resources. The goal of the surrogate model is to approximate the original model’s behavior, achieving accuracies close to the original model, at substantially reduced computational cost and time. In time-critical epidemic or pandemic situations, the evaluation of a large set of potential reactions is recommended, and swift reaction is crucial kuhn_regional_2022 . Unfortunately, supercomputing resources might not be available on model execution or predictions might be needed by stakeholders which do not have access to supercomputing facilities. To overcome these barriers and provide just-in-time answers to urgent questions, (pre)trained, e.g., overnight, surrogate models can become of great aid for decision makers and substantially increase the impact of mathematical-epidemiological modeling. With a regularly calibrated mechanistic model that integrates most recent viral parameter knowledge, mechanisms, and changed contact behavior, the surrogate model complements classical, purely data-driven ML models through the learning step of the simulator.
In this article, we present a surrogate model based on neural networks that is capable of handling temporal and spatial resolution. While several recent surrogate models are based on the concept of physics-informed neural networks calzolari_deep_2021 , sun_surrogate_2020 , shi_physics-informed_2022 , our surrogate model is entirely data-driven and uses a “physics”-based simulator only to create a wide range of reliable training data. On the one hand, this means that physical constraints such as population conservation are not considered explicitly. On the other hand, our approach offers a substantial advantage, as particular modifications to the original model do not require changes in the surrogate model architecture; only new training data has to be generated and the model has to be retrained, potentially using the previous model as a starting point. Since (epidemiological) parameters usually do not change on a daily basis, retraining is only necessary if these parameters change substantially. In future work, we will investigate if our surrogate model could further benefit from integrating “physical” constraints such as population conservation into the loss function.
@@
In this study, we first compare four different types of GNN layers using the binary adjacency matrix modeling the mobility between different spatial regions as input. Therefore, although already providing convincing results, none of the considered approaches takes the highly varying values for traveling and commuting activities between the different spatial regions into account. We expect that incorporating this information as edge weights in the input graph could further improve the performance. Promising directions for incorporating the edge weights could be the CrystalConv xie_crystal_2018 or XENetConv maguire_xenet_2021 layers. However, due to hardware limitations, specifically GPU memory on a single NVIDIA Quadro RTX 6000, we were unable to train models incorporating large graphs with 40,000 edges. Addressing this limitation will be a focus of future work.
This study mainly sheds light on the performance of surrogate model using the same graph connectivity in training and evaluation processes. We only briefly explored the generalization potential of the GNN model to other graphs than the one used for training, which is one of the strengths of GNNs. We considered the mobility network of 400 German counties and removed different quantities of edges, observing that a modest topological change with 20 % removed edges still yields acceptable prediction results. In particular, research on generalization with several realistic mobility networks and edge weights will be an avenue for future research.
@@
Our surrogate model delivers an age-structured and spatially resolved disease dynamics for a prediction horizon of several months and speeds up the prediction process substantially. As these results can be computed within a fraction of a second, our model is capable of being integrated in web-based application.
@@ Acknowledgements
All data and code used to generate the results presented in this work are publicly available as part of the MEmilio framework Bicker_MEmilio_2025 . The specific implementation used in this study, including all relevant scripts and pre-trained models, can be found under a persistent doi at zenodo schmidt_2025_15055792 .
@@
Funding
This work was supported by the Initiative and Networking Fund of the Helmholtz Association (grant agreement number KA1-Co-08, Project LOKI-Pandemics), by the German Federal Ministry for Digital and Transport (grant agreement FKZ19F2211A, Project PANDEMOS), and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) (grant agreement 528702961). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Competing Interests
All authors declare to not have any competing interests.
Appendix A Supplementary Figures